### v1

@@ 8.3 Related Literature
The third strand is the literature on the core. When modeling a matching problem under complete information, the core of the marriage market is the set of stable matchings (Roth and Sotomayor, 1990, Sasaki and Toda, 1992). For exchange economies with asymmetric information, Wilson (1978) introduced the notions of the coarse core and the fine core, which involves no information sharing or maximal information sharing respectively. Since both the core allocations and counterfactual blocking may be informative, these core concepts have been further refined (e.g., Vohra, 1999, Forges, 2004, Dutta and Vohra, 2005, Ray and Vohra, 2015). For surveys of the core under incomplete information, see Forges, Minelli, and Vohra (2002) and Forges and Serrano (2013). The coarse core has been extended to coalitional games with unawareness by Bryan, Ryall, and Schipper (2022). Our notion of stability is inspired by the coarse core. Our notion of flirt-proof stability is inspired by refinements of the coarse core that feature intentional communication. In some sense, flirt-proof stability combines the idea of the coarse core and with the idea of non-cooperative disclosure games (e.g., Milgrom and Roberts, 1986) but keeps the spirit of cooperative game theory.
@@
8.4 Possible Extensions
While we focus in this paper on unawareness and perfect information for simplicity, ordinality of preferences, and our desire to isolate phenomena related to unawareness independent from imperfect information, it would be interesting to allow for imperfect information as well. Players could be uncertain about the awareness of others (i.e., including the experiences they made in the past) and their preferences. Matchings and absence of blocking become informative. Moreover, both matchings and flirting can be tools for experimentation with the aim not just to learn about what one is imperfectly informed about but also discover what one is unaware. While it is straightforward to extend the unawareness structure to allow for imperfect information using unawareness type spaces introduced in Heifetz, Meier, and Schipper (2013), finding the appropriate solution concept is not straightforward. As it has been observed in the literature on the core under asymmetric information and matching with incomplete information, a satisfactory solution concept to matching under incomplete information should take into account the information revealed in matchings and counterfactual blocking (for surveys, see Forges and Serrano, 2013, Forges, Minelli, and Vohra, 2002). Even without unawareness, progress has been made almost exclusively only in matching games with one-sided incomplete information (e.g., Liu et al, 2014, Bikhchandani, 2017, Liu, 2020).
While our analysis is mainly cast in terms of the marriage market (again for simplicity and as a guide for intuitions), we believe that it can be extended to other matching problems including settings with transferable utility and many-to-many matching. Unanticipated experiences may happen in employee-employer relationships, school choice, with “experience” goods, etc. Similarly, within the current setting of the marriage market, alternative divorce rules may be explored. Currently, divorce occurs when a players unilaterally separates the match. Yet, human history also saw other divorce rules such as the requirement of bilateral agreement to separate the link. We conjecture that such an alternative divorce rule leads to less divorces, less experiences and thus less awareness within the market.
This paper focuses on decentralized matching under unawareness. It would be interesting to consider centralized matching in matching mechanisms. One desideratum for such mechanisms would be to pool awareness of all participants and implement an efficient matching at the pooled awareness level. Recently, Pram and Schipper (2025) introduced dynamic direct elaboration mechanisms with VCG-type transfers in quasi-linear environments to implement efficiently under pooled awareness. We currently consider extensions to matching.
Appendix A Application of the Cyclic Example to Two-Sided Market Entry
In this section, we present another application of our observation in Section 2.2 to market entry in matching markets. Bennett (1994) considers adding a single agent (say, a woman) to a marriage market with stable matching, and shows that the process (i.e., optimal divorce and remarriage procedure) of satisfying the optimal blocking pair for the unmatched woman terminates in a finite number of steps and reaches a stable matching. We consider an extended problem by adding both a woman and a man to marriage market with a stable matching and allow satisfying of mutually optimal blocking pairs. In contrast to Bennett’s problem of adding a single agent only, a modification of our example shows that this process can now lead to a cycle.
The initial matching market consists of three man $\{m_{1},m_{3},m_{4}\}$, and three woman $\{w_{2},w_{3},w_{4}\}$. Let the preference profile be consistent with the preferences in Example 2. That is, we consider the rank-order lists as in Example 2 but cross out $m_{2}$ and $w_{1}$:
[…]
Let the initial matching be $\mu_{1}=\begin{pmatrix}m_{1}&m_{3}&m_{4}\\ w_{2}&w_{3}&w_{4}\end{pmatrix}$, which is stable (without $m_{2}$ and $w_{1}$). Add $m_{2}$ and $w_{1}$ to the market. The unique mutually optimal blocking pair is $(m_{2},w_{2})$
Satisfying $(m_{2},w_{2})$ leads to the matching $\mu_{2}=\begin{pmatrix}m_{1}&m_{2}&m_{3}&m_{4}&w_{1}\\ m_{1}&w_{2}&w_{3}&w_{4}&w_{1}\end{pmatrix}$, which is the same as in the cycle in Example 2. As illustrated before, there is always a unique mutually optimal blocking pair along this path, and it cycles back to $\mu_{2}$. This observation shows that Bennett’s optimal divorce and remarriage procedure cannot easily be extended to situations where both a man and a woman are added to the market at the same time.

### v2

@@ 7.3 Related Literature
The third strand is the literature on the core. When modeling a matching problem under complete information, the core of the marriage market is the set of stable matchings (Roth and Sotomayor, 1990, Sasaki and Toda, 1992). For exchange economies with asymmetric information, Wilson (1978) introduced the notions of the coarse core and the fine core, which involves no information sharing or maximal information sharing respectively. Since both the core allocations and counterfactual blocking may be informative, these core concepts have been further refined (e.g., Vohra, 1999, Forges, 2004, Ray and Vohra, 2015). For surveys of the core under incomplete information, see Forges, Minelli, and Vohra (2002) and Forges and Serrano (2013). The coarse core has been extended to coalitional games with unawareness by Bryan, Ryall, and Schipper (2022). Our notion of stability is inspired by the coarse core. Our notion of flirt-proof stability is inspired by refinements of the coarse core that feature intentional communication. In some sense, flirt-proof stability combines the idea of the coarse core and with the idea of non-cooperative disclosure games (e.g., Milgrom and Roberts, 1986) but keeps the spirit of cooperative game theory.